OpenStudy (anonymous):

6 + x + y = x + 6 + y is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

the symbols moved (changed places) that is the distributive law

jimthompson5910 (jim_thompson5910):

no the distributive law would be x(y+z) = xy + xz

OpenStudy (anonymous):

If an operation is symmetrical (e.g. x+y=y+x) it is known as also being commutative.

OpenStudy (anonymous):

so is it a or c?

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

wow i really slept that one!!

OpenStudy (anonymous):

symbols change places it is COMMUTATIVE law

OpenStudy (anonymous):

Jaja, it's cool, mate. Also, recall that distributive property requires two distinct operations on a field.